OpenStudy (anonymous):

Help please!!!!!! Giving medals! Write x2 - 8x + 13 = 0 in the form (x - a)2 = b, where a and b are integers. (x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4

OpenStudy (cggurumanjunath):

a

OpenStudy (cggurumanjunath):

(x - 4)^2 = 3

OpenStudy (anonymous):

Thank you :) do you mind helping me with another question?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

sure .

OpenStudy (anonymous):

Ok thanks :) The function f(x) = -x^2 - 9x + 10 shows the relationship between the vertical distance of a diver from a pool's surface f(x), in feet, and the horizontal distance x, in feet, of a diver from the diving board. What is a zero of f(x) and what does it represent? x = 1; the diver jumps in the pool at 1 foot per second x = 10; the diver jumps in the pool at 10 feet per second x = 1; the diver hits the water 1 foot away horizontally from the board x = 10; the diver hits the water 10 feet away horizontally from the board

OpenStudy (cggurumanjunath):

zero of a function is f(0) =10

OpenStudy (anonymous):

So its either B or D?

OpenStudy (cggurumanjunath):

ya.

OpenStudy (cggurumanjunath):

D

OpenStudy (anonymous):

Alright, thanks! Can you help me with one more?

OpenStudy (cggurumanjunath):

SURE.GO AHEAD.

OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 - x - 2? x = -2, 1 x = 2, -1 x = -2, -1 x = 2, 1 A polynomial function can be written as (x + 1)(x + 4)(x - 7). What are the x-intercepts of the graph of this function? (1, 0), (4, 0), (7, 0) (-1, 0), (-4, 0), (-7, 0) (1, 0), (4, 0), (-7, 0) (-1, 0), (-4, 0), (7, 0)

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

(-1, 0), (-4, 0), (7, 0)

OpenStudy (cggurumanjunath):

http://prntscr.com/2mcx7e for previous one.